OpenStudy (anonymous):

Can someone please help me solve this function? Given f(x)=5x-8 and g(x)=x^2+4 Find (f+g)(2) and (f/g)(x)

OpenStudy (anonymous):

(f+g)(x)=5x-8 +x^2+4=x^2+5x-4 (f+g)(2)=2^2+5*2-4=4+10-4=10

OpenStudy (anonymous):

(f/g)(x)=( 5x-8)/( x^2+4)